The position of a squirrel running in a park is given by

At 5.95 s, how far is the squirrel from its initial position and what is the magnitude of the squirrel's velocity?
Answer #1
after 5.95 sec
r = 2.94i + 4j
initial position =0i +0j

displacement = (2.94^2+4^2)^0.5 = 4.966m
velocity = dr / dt = (0.28+0.072t)i + 0.057t^2j
at time t = 5.95 , Vt = 0.7084i + 2.0179j m/sec
Answer #2
[REMOVED]